OpenStudy (anonymous):

The perimeter of a rectangle is 68 feet and its width is 8/9 times its length. Find the dimensions of the rectangle

jigglypuff314 (jigglypuff314):

Perimeter of rectangle = 2(length) + 2(width) so given P = 68 and L = 8/9W then 68 = 2(8/9 W) + 2W 68 = 16/9 W + 18/9 W 68 = 34/9 W W = 18 then plug back into P = 2L + 2W so 68 = 2L + 2(18) then find L therefore W = ___ and L = ___

OpenStudy (anonymous):

Thank you so much!!

jigglypuff314 (jigglypuff314):

you're welcome ;)